# Coherent coupling of YBCO superconducting resonators and 100-nm-thick YIG films

### v1

@@ Abstract
In cavity magnonics, magnon-photon hybridization has been widely investigated for both fundamental studies and applications. Planar superconducting resonators operating at microwave frequencies have demonstrated the possibility to achieve high couplings with magnons by exploiting the confinement of the microwave field in a reduced volume. Here we report a study of the coupling of high-$T_{c}$ YBCO superconducting waveguides with 100-nm-thick YIG magnetic films. We study the evolution of mode frequencies at different temperatures and extract the coupling strength of hybrid magnon-photon modes. We show that the experimental results can be reproduced using a simple model in which the temperature dependence of the penetration depth accounts for the evolution of the polaritonic spectrum.
@@
Magnons , Spin waves , Microwaves , Strong coupling , Hybrid systems , Superconducting resonators
@@ 1 Introduction
Magnon-photon hybrids are obtained by coherently coupling collective excitations in coupled spin systems, that is spin waves and their quantized counterpart magnons, with microwave photon modes. These systems have proven to be fertile ground for both fundamental studies and applications. In the former case, topics like quantum magnonics [1, 2], nonlinear effects [3, 4, 5], ultrastrong coupling [6, 7, 8], non-Hermitian physics [9, 10], Floquet engineering [11], magnonic frequency combs [12] have been recently investigated. Applications in magnonics and quantum technologies, either at room temperature or in the quantum regime at mK temperatures, are related, but not restricted, to computation, sensing, microwave-to-optical transduction, coherent coupling of remote physical systems, emission of microwaves, among others [13, 14, 15, 16, 17].
@@
For these studies, magnetic materials such as electrically insulating yttrium iron garnet (YIG) or conducting permalloy, which are characterized by high spin densities and low ferromagnetic resonance (FMR) damping rates, have found widespread attention [18]. The incorporation of these materials into superconducting circuits has opened additional possibilities [19, 20, 21, 22]. Microwave modes in planar superconducting resonators have been exploited to achieve strong [23, 24, 25, 26] and ultrastrong [27, 28, 29, 8, 30] coupling with magnons, also using nanomagnets fabricated directly onto the resonator [25, 26]. Furthermore, superconducting qubits were coupled to YIG spheres to detect single-magnon excitations [1, 31].
@@
High critical temperature ($T_{c}$) superconducting resonators, in particular YBa2Cu3O7 (YBCO), show $T_{c}$ up to about 90 K and resilience in applied magnetic field [32]. Strong and ultrastrong coupling with spin systems have been achieved, respectively, using paramagnetic ensembles [33, 34, 35, 36] and 5-$\mu\mathrm{m}$-thick YIG films [8, 30]. In the latter case, the evolution of polaritonic modes at temperature below $T_{c}$ shows a progressive frequency shift of the hybrid magnon-photon mode, which has been attributed to the effect of Meissner currents and correlated with the penetration depth in the superconducting layer [30].
@@
Figure 1: (a) Sketch of the experimental geometry. (b) Spectral map showing the numerical derivative of broadband transmission spectra acquired at $T=30~{}\mathrm{K}$. Red and green dashed lines show $\omega_{0}$ and $\omega_{1}$ as derived respectively from Eq. 2 and Eq 4. The blue dashed line show $\omega_{FMR}$. (c) Evolution, as a function of the temperature, of the $S_{21}$ spectra measured at $\mu_{0}H_{0}=0.23~{}\mathrm{T}$. The incident microwave power is $-15~{}\mathrm{dBm}$. The background transmission, derived from high magnetic field data, was subtracted from the spectra. Solid lines display the Lorentzian fit (Eq. 1).
@@
The spectrum of spin-wave excitations in the YIG film and the coupling with the resonator are dependent upon the thickness of the magnetic layer. We have experimentally investigated the interplay between YBCO and 100-$\mathrm{nm}$-thick YIG films. We first study the broadband spectrum of excitations acquired using a YBCO coplanar waveguide (CPW) transmission line at different temperatures. We then analyze the coupling between microwave and magnon modes using a CPW resonator. We finally discuss how the extracted physical quantities evolve by varying the thickness of the YIG film.
@@ 2 Broadband transmission spectroscopy
We studied a 100-nm-thick YIG film that was grown by liquid phase epitaxy on a 0.5-mm-thick Gadolinium Gallium Garnet (GGG) substrate. The magnetic film was positioned above the YBCO layer and gently pushed from the backside. Transmission ($S_{21}$) spectra were measured as a function of frequency using a broadband CPW line with the external magnetic field ($H_{0}$) applied parallel to the central conductor (Fig. 1(a)). At low temperature ($T=30~{}\mathrm{K}$), the transmission derivative ($\partial S_{21}/\partial H_{0}$) map evidences the presence of a main resonance mode at the lowest frequency ($\omega_{0}$) plus an additional mode ($\omega_{1}$), which has much weaker amplitude, shifted by approximately $1.2~{}\mathrm{GHz}$ with respect to the former (Fig. 1(b)).
As a function of frequency, transmission spectra taken at $\mu_{0}H_{0}=0.23~{}\mathrm{T}$ show well-defined absorption dips (Fig. 1(c)). As the temperature decreases, the resonance frequency, $\omega_{0}$, progressively moves toward higher frequencies. The spectral lineshape of the absorption dip, obtained by subtracting the background from the transmission spectra, can be fitted using a Lorentzian curve [37]
@@
where $\omega$ is the frequency, $a$ is the amplitude and $\Delta\omega$ is the full width at half maximum linewidth (Fig. 1(c)). In particular, from the fits we note that $\Delta\omega$ increases from $30~{}\mathrm{MHz}$ at 80 K to $64~{}\mathrm{MHz}$ at 10 K. These values are influenced by the presence of both paramagnetic GGG substrate and superconducting YBCO layer as we discuss in the following.
Fig. 2 shows broadband spectral maps taken at temperatures between 80 and 10 K. In order to make the position of the YIG modes more evident, we plot the second derivative of the transmission with respect to the magnetic field $\partial^{2}S_{21}/\partial H_{0}^{2}$. We note that spectra taken above 50 K show the presence of $\omega_{0}$ only, while $\omega_{1}$ is visible at 30 K and below. $\omega_{1}(H_{0})$ curve varies only slightly from 10 K to 30 K (Fig. 2).
Figure 2: Evolution of broadband spectral maps measured at different temperatures. The color scale shows the second derivative of transmission with respect to the magnetic field. Dashed lines show the curves calculated with Eq. 2 and Eq. 4, as indicated.
Considering the lateral profile of the CPW broadband line and the Damon-Eshbach geometry in our experiment, we expect that the transmission line excites spin wave modes with wavenumber up to $k_{y}=3\times 10^{5}~{}\mathrm{rad~{}m}^{-1}\approx 2\pi/s$ [38, 30]. According to the analytical model by Kalinikos ans Slavin (KS) [39], the frequency of the lowest mode can be calculated as [40]
@@
where $\gamma=28.02~{}\mathrm{GHz/T}$ is the electron gyromagnetic ratio and $P_{00}=1+[(1-\exp(-k_{y}d))/k_{y}d]$. We note that, being $d=100~{}\mathrm{nm}$ and $k_{y}d=0.03$, Eq. 2 results very near to $\omega_{FMR}=\mu_{0}\gamma\sqrt{H_{0}(H_{0}+M_{s})}$ (Fig. 1(a)). The temperature dependence of Eq. 2 derives from the saturation magnetization of YIG, $M_{s}$, which follows [41, 42, 30]
@@
being $u=23\times 10^{-6}~{}\mathrm{K^{-3/2}}$ and $v=1.08\times 10^{-7}~{}\mathrm{K^{-5/2}}$ [42]. Using $\mu_{0}M_{0}=0.28$ T, we can reproduce the magnetic field dependence of $\omega_{0}$ in the entire temperature range (Fig. 1(a) and Fig. 2).
@@
where $A$ is the exchange constant of the YIG film. The experimental spectra in Figs. 1 and 2 can be reproduced by using $A=5.2~{}\mathrm{pJ/m}$ at 30 K and $A=5.5~{}\mathrm{pJ/m}$ at 10 K.
@@ 3 Coupling between magnetic film and resonator
The 100 nm thick YIG film was subsequently placed on top of the YBCO resonator to test the coupling between their modes. Transmission spectra were acquired between 10 K and 85 K with the film in contact with the resonator. Fig. 3 shows the formation of two polaritonic branches, indicating hybridization between the magnon and photon modes. At low temperature, the splitting of approximately 460 MHz corresponds to the collective coupling $g\approx 230~{}\mathrm{MHz}$. As the temperature increases up to the critical temperature of YBCO, panels (a-f) show a progressive lowering of the resonator frequency ($\omega_{c}$) along with a decrease in polariton splitting.
Figure 3: Transmission ($\partial S_{21}/\partial H_{0}$) spectra acquired at different temperatures. The YIG film is positioned in contact with the resonator and the incident microwave power is $-15~{}\mathrm{dBm}$. For comparison, spectra taken in the same conditions with the GGG face of the sample located above the resonator do not show any resonant mode. The calculated polaritonic modes $\Omega_{\pm}$ are represented by red dashed lines; blue and yellow dashed lines show $\omega_{0}$ and $\omega_{b}$, respectively.
@@
Figure 4: Frequency of the resonator extracted from transmission maps acquired at different temperatures (circles). The solid line shows the fit with Eq. 9. Inset. Coupling strength and frequency shift calculated using, respectively, the resonator frequencies (circles) and Eq. 9 (solid lines).
@@
being $L[\lambda_{L}(T)]$ the temperature-dependent inductance and $\omega_{c}(0)=9.7~{}\mathrm{GHz}$ at $T_{0}=10~{}\mathrm{K}$. Finally, the collective coupling $g=g_{s}\sqrt{2s_{\mathrm{Fe}}N_{s}}$, where $s_{\mathrm{Fe}}=5/2$ is the single-ion spin of Fe3+, is calculated assuming that the number of spins is $N_{s}=1.49\times 10^{13}$ in the entire temperature range. The temperature dependence of $g$ derives, through $\omega_{c}(T)$, from the spin-photon coupling
@@
In summary, Fig. 4 shows the temperature evolution of the parameters that describe the experimental data. From the fitted values of $\omega_{c}(T)$ and using the values of $\delta_{sc}(T)$ and $g(T)$ calculated using $r=0.45$ and $N_{s}=1.49\times 10^{13}$, we can reproduce the evolution of the polaritonic spectrum in the 10-85 K range (Fig. 3).
@@ 4 Discussion
Data acquired at low temperature using the YBCO broadband line (Fig. 1(b) and Fig. 2) evidence the presence of two magnetic modes having frequencies $\omega_{0}$ and $\omega_{1}$. Their evolution as a function of magnetic field and temperature has been analyzed using the KS model and taking into account the saturation magnetization ($M_{s}$) and exchange constant ($A$). The values found for such parameters (Sect. 2) are in line with those reported in the literature [42, 46]. However, because the magnetic film is in contact with the surface of the superconductor, additional effects emerge due to their interplay [30] as evidenced by the comparison of transmission spectra taken at different temperatures.
The evolution of broadband maps (Fig. 2) shows that $\omega_{1}$ becomes visible only at $T\leq 30~{}\mathrm{K}$. The saturation magnetization and exchange constant of YIG have a weak temperature dependence in this range; by contrast, the characteristics of the YBCO film are strongly temperature dependent. In particular, the penetration depth decreases at low $T$, where the microwave field becomes more efficiently confined to the surface of the transmission line. In these conditions, similar to [8, 30], we note that the transmission line becomes more sensitive to spin wave modes having a short decay distance of the stray field, such as for high wavenumbers [38, 47].
The linewidths extracted from the broadband spectra in Fig. 1(c) show values larger than what is typically reported for YIG/GGG films. Losses in YIG/GGG films have been widely investigated as a function of experimental parameters such as temperature, film thickness, and wavenumber [48, 49, 50]. At room temperature, FMR resonances show linewidths well below 1 MHz; at low temperature the broadening of spin wave modes has been attributed to effects like dipolar coupling to the partially magnetized GGG substrate and relaxation of rare earth ions, resulting in linewidths between 1 and 30 MHz. Moreover, studies of magnetic losses in YIG films having conductive electrodes deposited on top have shown increased values of linewidth [51, 52, 53]. In our case, the wide linewidths in Fig. 1(c) can be attributed to the presence of the superconductor. We just mention that, due to Meissner screening and nonperfect parallelism of the film surface with respect to magnetic field, we expect field inhomogeneity at the superconductor-magnetic film interface.
Further evidence of the interplay between magnetic and superconducting layers emerges in transmission spectra acquired using the superconducting resonator. In the first place, we note that $\omega_{1}$ is not visible in Fig. 3, therefore, we can rule out the effects of higher modes observed with three-dimensional cavities [54]. Following the results obtained with 5-$\mathrm{\mu m}$-thick YIG films [30], in Sect. 3 we have assumed that $\omega_{b}$ is derived by the shift of the lowest YIG mode, $\omega_{0}$ (Eq. 6), as quantified by Eq. 7. In particular, the latter predicts that $\delta_{sc}$ is proportional to the thickness of the YIG film ($d$): the maximum value of $\delta_{sc}$ obtained in Ref. [30] for $d=5~{}\mathrm{\mu m}$ exceeds $1~{}\mathrm{GHz}$ whereas in the present case ($d=100~{}\mathrm{nm}$) $\delta_{sc}$ drops to $0.1~{}\mathrm{GHz}$ (Fig. 4). However, we notice that the geometrical factor $r$ used to fit the spectra is different in the two cases. These parameters well reproduce the evolution of the transmission data in Fig. 3. The absence of a noticeable frequency shift in the broadband spectra (Fig. 2) could be related to the relatively wide linewidth of $\omega_{0}$ (Fig. 1(c)).
It is interesting to compare the coupling strength values ($g$) obtained with YIG films having different thicknesses, though similar lateral dimensions, and obtained using YBCO resonators having the same geometry. The maximum coupling strength derived from the spectra in Fig. 3 is $0.23~{}\mathrm{GHz}$ while the coupling obtained with a 5-$\mathrm{\mu m}$-thick YIG film is around $1.1~{}\mathrm{GHz}$ [30]. We first note that the ratio between the thicknesses is 50 while the coupling strength achieved with the thicker film is 5 times higher; considering that $g\propto\sqrt{N_{s}}$, where $N_{s}$ is the product of the YIG spin density and the intersection of the resonator and YIG film volumes, we infer that the coupling strength does not scale as the square root of the thickness of the latter. In addition, the distribution of the microwave field obtained from the results of electromagnetic simulations, at least for the bare CPW resonator, stretches well above a few microns from the resonator surface [8]. If the magnetic system was paramagnetic or if the mode had uniform precession of the magnetization, all the spins, positioned in a 100-nm-thick layer above the surface of the resonator and located sufficiently near to the central conductor, were uniformly excited by the microwave field. However, the coupling strength estimated using the spin density of the YIG film results much higher than what derived from the experiment; this indicates that in the latter case the precession of the spins is nonuniform, i.e. the spin texture of the hybrid mode is more complex. Thus, while the spin-photon coupling is determined by the vacuum magnetic field of the resonator (Eq. 10), the collective coupling, as suggested by simple calculations, is related to the specific way in which the electromagnetic and spin waves combine in the magnet/superconductor bilayer to shape the hybrid mode.

### v2

# Coherent coupling between YBCO superconducting resonators and sub-micrometer-thick YIG films

@@ Abstract
In cavity magnonics, magnon-photon hybridization has been widely investigated for both fundamental studies and applications. Planar superconducting resonators operating at microwave frequencies have demonstrated the possibility to achieve high couplings with magnons by exploiting the confinement of the microwave field in a reduced volume. Here we report a study of the coupling of high-$T_{c}$ YBCO superconducting waveguides with 104-nm-thick YIG magnetic films. We study the evolution of mode frequencies as a function of temperature and extract the coupling strength of hybrid magnon-photon modes. We show that the experimental results can be reproduced using a simple model in which the temperature dependence of the penetration depth accounts for the evolution of the polaritonic spectrum.
@@
YIG , YBCO , Magnons , Microwaves , Strong coupling , Hybrid systems , Superconducting resonators
@@ 1 Introduction
Magnon-photon hybrids are obtained by coherently coupling collective excitations in coupled spin systems, that is spin waves and their quantized counterpart magnons, with microwave photon modes. These systems have proven to be fertile ground for both fundamental studies and applications [1]. In the former case, topics like quantum magnonics [2, 3], nonlinear effects [4, 5, 6], ultrastrong coupling [7, 8, 9], non-Hermitian physics [10, 11], Floquet engineering [12], magnonic frequency combs [13] have been recently investigated. Applications in magnonics and quantum technologies, either at room temperature or in the quantum regime at mK temperatures, are related, but not restricted, to computation, sensing, microwave-to-optical transduction, coherent coupling of remote physical systems, and emission of microwaves, among others [14, 15, 16, 1, 17].
@@
For these studies, magnetic materials such as electrically insulating Yttrium Iron Garnet (YIG) or conducting permalloy, which are characterized by high spin densities and low ferromagnetic resonance (FMR) damping rates, have found widespread attention [18]. Incorporation of these materials into superconducting circuits has opened additional possibilities [19, 20, 21, 22]. Microwave modes in planar superconducting resonators have been exploited to achieve strong [23, 24, 25, 26] and ultrastrong [27, 28, 29, 9, 30] coupling with magnons, also using nanomagnets fabricated directly onto the resonator [25, 26]. Furthermore, superconducting qubits were coupled to YIG spheres to detect single-magnon excitations [2, 31].
@@
High critical temperature ($T_{c}$) superconducting resonators, in particular YBa2Cu3O7 (YBCO), show $T_{c}$ up to about 90 K and resilience in applied magnetic field [32]. Strong and ultrastrong coupling with spin systems have been achieved, respectively, using paramagnetic ensembles [33, 34, 35, 36, 37] and few-$\mu\mathrm{m}$-thick YIG films [9, 30]. In the latter case, the evolution of polaritonic modes at temperature below $T_{c}$ has shown a progressive frequency shift of the hybrid magnon-photon mode, which has been correlated with the penetration depth in the superconducting layer as an effect of the interplay between Meissner currents and spin waves [30].
@@
Figure 1: Top view (a) and vertical section (b) of the CPW broadband line with the YIG film placed in top. Dashes in (b) show the distribution of the oscillating magnetic field $\mathbf{h_{ac}}$. (c) Evolution, at different temperatures, of the transmission spectra measured at $\mu_{0}H_{0}=0.23~\mathrm{T}$. The background $S_{21}$ spectrum obtained from off resonance data has been subtracted. Solid lines display the Lorentzian fit (Eq. 1). (d) Spectral map showing the numerical derivative of broadband transmission spectra acquired at $T=30~\mathrm{K}$. Green and yellow dashed lines show $\omega_{0}$ and $\omega_{1}$ as derived respectively from Eq. 2 and Eq 4. The red dashed line show $\omega_{FMR}$.
@@
The spectrum of spin-wave excitations in the YIG film and the coupling with the resonator depend upon the thickness of the magnetic layer. Here we experimentally investigate the evolution of transmission spectra obtained by placing a 104-$\mathrm{nm}$-thick YIG film both on a coplanar waveguide (CPW) broadband line and half-wavelength resonator fabricated from a YBCO film. We first study the broadband spectrum of excitations acquired using the YBCO CPW at different temperatures. We then analyze the coupling between microwave and magnon modes using the CPW resonator. We finally discuss how the extracted physical quantities compare with those obtained with thicker YIG films, which have been previously reported in Refs. [9, 30].
2 Experimental methods
CPW broadband line and resonator were fabricated from a YBCO film having thickness of $330~\mathrm{nm}~(\pm 10\%)$, which was deposited by reactive co-evaporation (RCE) [38] on a 3”-wide, 0.43-$\mathrm{mm}$-thick, r-cut sapphire substrate with a 20-nm-thick CeO2 buffer layer (Ceraco GmbH, M-type). The characteristic porous surface of the YBCO film shows enhanced flux pinning and critical current density higher than $2\times 10^{6}~\mathrm{A/cm^{2}}$ at 77 K. After deposition, the substrate was diced into $8\times 5\times 0.43~\mathrm{mm^{3}}$ blocks, which were patterned using optical lithography and dry etching by Ar plasma in a Reactive Ion Etching (RIE) chamber. The central conductor of the CPW line has characteristic width $w=17~\mathrm{\mu m}$ and separation $s=14~\mathrm{\mu m}$ between the central conductor and the lateral ground planes (Fig. 1(a,b)). The broadband CPW supports transmission up to 14 GHz; the half-wavelength resonator has the same geometry as the CPW broadband line except for two 140-$\mathrm{\mu m}$-wide input and output capacitive gaps that interrupt the central conductor to define a central strip having length of 6 mm [9, 30].
The YIG film that was grown by liquid phase epitaxy with thickness of 104 nm on a 0.5-mm-thick Gadolinium Gallium Garnet (GGG) substrate with (111) orientation (Matesy GmbH) [39]. The reported FMR linewidth is less than 2 Oe at room temperature.
The YIG/GGG film was cut in a $\approx 4\times 2~\mathrm{mm^{2}}$ sample and placed on the superconducting CPW. The film is gently pushed in contact with the YBCO surface by a plastic screw from the GGG side. The device was cooled down in a cryogenic setup having a superconducting solenoid that generates the external magnetic field ($H_{0}$), which is applied in the YIG film plane along the central conductor of the CPW (Fig. 1(a,b)). Transmission ($S_{21}$) measurements were carried out using a Vector Network Analyzer by sweeping the frequency at stationary values of the magnetic field. The incident microwave power is $\approx-8~\mathrm{dBm}$ at the CPW input port. Test experiments carried out in the same conditions using a bare GGG sample have not shown any detectable magnetic resonance lines. The derivatives of the transmission, $\partial S_{21}/\partial H_{0}$ and $\partial^{2}S_{21}/\partial H_{0}^{2}$, were calculated numerically.
@@ 3 Broadband transmission spectroscopy
We first used the broadband superconducting waveguide to study the evolution of ferromagnetic resonance lines. Transmission spectra evidence the presence of a main resonance at the frequency $\omega_{0}$ (Fig. 1(c,d)). As the temperature ($T$) decreases, $\omega_{0}$ progressively shifts toward higher frequencies. This trend indicates a decrease in the fixed-frequency resonance field as the temperature decreases between 80 K and 30 K. We fit the absorption dip, obtained by subtracting the background from the transmission spectra, using a Lorentzian curve [40]
@@
where $\omega$ is the frequency, $a$ is the amplitude and $\Delta\omega$ is the full width at half maximum (Fig. 1(c)). We note that $\omega_{0}/2\pi$ varies between 9.58 GHz at 80 K to 9.64 GHz at 10 K. Although the Lorentzian line shape does not perfectly reproduce the resonance lines in Fig. 1(c) we can obtain an estimate of $\Delta\omega/2\pi$, which increases from $30~\mathrm{MHz}$ at 80 K to $64~\mathrm{MHz}$ at 10 K.
When the YIG and YBCO layers are in good contact, an additional faint line can be observed at low temperature, whose frequency $\omega_{1}/2\pi$, is approximately $1.2~\mathrm{GHz}$ above $\omega_{0}/2\pi$ (Fig. 1(d)). Fig. 2 shows transmission maps taken at temperatures between 80 and 10 K, plotted as second derivative $\partial^{2}S_{21}/\partial H_{0}^{2}$ to evidence $\omega_{0}$ and $\omega_{1}$. We note that spectra taken above 50 K show the presence of $\omega_{0}$ only, while $\omega_{1}$ is visible at 30 K and below.
Figure 2: Evolution of spectral maps measured at different temperatures with the broadband CPW. The color scale shows the second derivative of transmission with respect to the magnetic field. Dashed lines display the curves calculated with Eq. 2 and Eq. 4, as indicated.
Considering the lateral profile of the broadband CPW and the Damon-Eshbach geometry in our experiment, we expect that the transmission line excites spin wave modes with wavenumber up to $k_{y}=3\times 10^{5}~\mathrm{rad~m}^{-1}\approx 2\pi/s$ [41, 30]. According to the analytical model by Kalinikos ans Slavin (KS) [42], the frequency of the lowest mode can be calculated as [43]
@@
where $\gamma=28.02~\mathrm{GHz/T}$ is the electron gyromagnetic ratio and $P_{00}=1+[(1-\exp(-k_{y}d))/k_{y}d]$. We note that, being $d=104~\mathrm{nm}$ and $k_{y}d=0.03$, Eq. 2 results very close to $\omega_{FMR}=\mu_{0}\gamma\sqrt{H_{0}(H_{0}+M_{s})}$ (Fig. 1(d)). The temperature dependence of Eq. 2 derives from the saturation magnetization of YIG, $M_{s}$, which follows [44, 45, 30]
@@
being $u=23\times 10^{-6}~\mathrm{K^{-3/2}}$ and $v=1.08\times 10^{-7}~\mathrm{K^{-5/2}}$ [45]. Using $\mu_{0}M_{0}=0.28$ T, we can reproduce the magnetic field dependence of $\omega_{0}$ in the entire temperature range (Fig. 1(c,d) and 2).
@@
where $A$ is the exchange constant of the YIG film. The experimental spectra in Figs. 1 and 2 can be reproduced using $A=5.2~\mathrm{pJ/m}$ at 30 K and $A=5.5~\mathrm{pJ/m}$ at 10 K.
@@ 4 Coupling between magnetic film and resonator
The 104-nm-thick YIG film was subsequently placed on top of the superconducting resonator to test the coupling between magnon and photon modes. At low temperature, the bare YBCO/sapphire resonator shows the fundamental mode at $\omega_{c}/2\pi\approx 10~\mathrm{GHz}$, which decreases to $\omega_{c}/2\pi\approx 9.7~\mathrm{GHz}$ in the presence of the YIG/GGG sample.
Magnon-photon hybridization is shown by the formation of polariton branches in transmission spectra displayed for a series of temperatures in Fig. 3. Below 50 K (panels (a-c)), the splitting of approximately 460 MHz corresponds to the collective coupling $g/2\pi\approx 230~\mathrm{MHz}$. As the temperature increases, we observe the progressive lowering of the resonator frequency and the decrease in polariton splitting (panels (d-f)). Above $T_{c}$ the anticrossing disappears.
Figure 3: Evolution of transmission ($\partial S_{21}/\partial H_{0}$) spectra acquired as a function of temperature using the CPW resonator
. The calculated polaritonic modes $\Omega_{\pm}$ are represented by red dashed lines; blue and yellow dashed lines show $\omega_{0}$ and $\omega_{b}$, respectively.
@@
Figure 4: Frequency of the resonator extracted from transmission maps acquired at different temperatures (circles). The solid line shows the fit with Eq. 9. Inset. Coupling strength and frequency shift calculated using, respectively, the resonator frequencies (circles) and Eq. 9 (solid lines). Error bars are estimated from the uncertainty of the fit parameters.
@@
being $L[\lambda_{L}(T)]$ the temperature-dependent inductance and $\omega_{c}(0)/2\pi=9.7~\mathrm{GHz}$ at $T_{0}=10~\mathrm{K}$. Finally, the collective coupling $g=g_{s}\sqrt{2s_{\mathrm{Fe}}N_{s}}$, where $s_{\mathrm{Fe}}=5/2$ is the single-ion spin of Fe3+, is calculated assuming that the number of spins is $N_{s}=1.49\times 10^{13}$ in the entire temperature range. The temperature dependence of $g$ derives, through $\omega_{c}(T)$, from the spin-photon coupling
@@
Fig. 4 shows the temperature evolution of the parameters that describe the experimental data. From the fitted values of $\omega_{c}(T)$ and using the values of $\delta_{sc}(T)$ and $g(T)$ calculated with $r=0.45$ and $N_{s}=1.49\times 10^{13}$, we can reproduce the evolution of the transmission spectra shown in Fig. 3 for temperatures between 10 and 85 K.
@@ 5 Discussion
It is worth comparing the results obtained here for the 104-nm-thick YIG film with those reported for thicker films [9, 30]. The broadband spectra shown in Figs. 1 and 2 display two well-separated resonances at frequencies $\omega_{0}$ and $\omega_{1}$, in contrast with the larger number of closely spaced lines observed for a 5-$\mathrm{\mu m}$-thick YIG film under similar experimental conditions [9, 30]. Their dependence as a function of $H_{0}$ has been analyzed with Eqs. 2 and 4, using fitting parameters compatible with those previously reported [45, 49, 9, 30]. In particular, Eq. 4 shows, in accordance with the trend observed in the experiments, an increase in the frequency $\omega_{1}$ when the thickness of the YIG film decreases. The chosen fitting parameters account for the non-trivial evolution of the FMR line at low temperature, which has recently been attributed to the increase in magnetic anisotropy as a result of the interaction between the YIG film and the GGG substrate [50, 51, 39]. We note that the temperature evolution of the resonance position in Figs. 1(c) and 2, and the deviation from the Lorentzian line shape in the resonance spectra, are in line with other recently reported experimental [50, 39].
The broadening of the lowest resonant peak and the diminishing of its amplitude at low temperature evidence the increase of losses (Fig. 1(c)). Despite the narrow FMR linewidths obtained at room temperature [52, 53], low-temperature measurements on YIG/GGG films have shown FMR linewidths up to 30 MHz due to local defects and other effects induced by the GGG substrate [50, 51, 39]. Moreover, experiments with YIG films having conductive electrodes on top have shown the broadening of the FMR line [54, 55, 56]. Additional contributions to line broadening come from the superconductor. Inhomogeneities in the local magnetic field can be related to Meissner screening and non-perfect parallelism of the film surface with respect to the external magnetic field. Being our YBCO film a type-II superconductor with pinned vortices, in the mixed state we also expect the presence of inhomogeneous magnetic fields at the interface with the YIG film due to the presence of the Abrikosov lattice.
Additional fingerprints of the interplay between magnetic and superconducting layers can be found in the temperature-dependent properties of the YBCO CPWs. The amplitude of the microwave field decreases quasiexponentially as the height from the CPW increases [9], additionally, at low $T$ the penetration depth of YBCO drops to $\approx 150$ nm [30]. The appearance of higher magnon modes in broadband measurements (Fig. 2) is likely due to the coupling between tightly confined microwaves and stray fields [41]. Furthermore, experiments with thick YIG films and YBCO resonators in the ultrastrong coupling regime evidenced the shift of the anticrossing position with respect to the unperturbed resonance line of the YIG film ($\delta_{sc}/2\pi$ up to $1.2~\mathrm{GHz}$) [30]. This behavior has been interpreted as an effect of electromagnetic proximity between the magnetic and superconducting layers [21, 57].
In the present case, the 104-nm-thick YIG film gives rise to a much smaller shift ($\delta_{sc}/2\pi=0.1~\mathrm{GHz}$), which is justified by the smaller $d$ in Eq. 7. However, the broadband spectra in Fig. 2 suggest that the line position at different temperatures can be reproduced using only the parameters of the YIG, without including the effect of the superconductor. We note that in Eq. 7 the geometric parameter $r$ has been used as a free fitting parameter to adjust the absolute value of $\delta_{sc}$; to better quantify its magnitude, a more systematic study that involves variations in the size and thickness of the YIG film would be required. Furthermore, we note that the estimated magnon and photon numbers are $N_{m}=N_{p}\approx 10^{10}\ll N_{s}$ [9]. The shifts in magnon frequency extracted from Fig. 3 are greater than those expected from the magnon Kerr effect in our experiment [58, 59]. Considering the estimated mode volume of the resonator $V_{m}=N_{s}/\rho=7\times 10^{-14}~\mathrm{m^{3}}$, being $\rho=2.1\times 10^{28}~\mathrm{m^{-3}}$ the spin density of the YIG, and the first-order anisotropy constant of the YIG film ($K_{an}$), we obtain the Kerr coefficient $K=\hbar\mu_{0}K_{an}\gamma^{2}/(M_{s}^{2}V_{m})\approx 10^{-7}~\mathrm{Hz}$. The Kerr effect results much lower than the collective spin-photon coupling. These numbers further confirm our analysis carried out in the linear regime with a low number of excitations.
We finally make a direct comparison between the coupling strength obtained and those reported in [9, 30] using thicker YIG/GGG films and identical YBCO CPW resonator under similar experimental conditions. The maximum coupling strength derived from the spectra in Fig. 3 is $0.23~\mathrm{GHz}$ while the coupling obtained with a 5-$\mathrm{\mu m}$-thick YIG film having similar lateral size is around $1.1~\mathrm{GHz}$ [30]. The ratio between thicknesses is 50 while the coupling strength achieved with the thicker film is only 5 times higher. By using wider 5-$\mathrm{\mu m}$-thick films, the coupling strength increased to around 2 GHz, although no significant improvement was observed for thicknesses as high as 20 $\mathrm{\mu m}$ [9]. These results confirm that in our experiments a significant coupling with the CPW is achieved at distances below $1~\mathrm{\mu m}$ above the YBCO surface.